OpenStudy (anonymous):

Write the prime factorization, using "x" between the numbers: 102

OpenStudy (anonymous):

102/2=51 51/3=17 17/17=1 2x3x17=102

OpenStudy (jamesj):

102 = 2x51 = 2x3x17